A proton accelerates from rest in a uniform electric field of 590 N/C. At one later moment, its speed is 1.30 Mm/s (nonrelativistic because v is much less than the speed of light).

(a) Find the acceleration of the proton.
____________ m/s2

(b) Over what time interval does the proton reach this speed?
_____________ s

(c) How far does it move in this time interval?
______________ m

(d) What is its kinetic energy at the end of this interval?
_______________ J
Answer #1

(a.) F = ma = Eq

so acceleration a = Eq / m = (590 * 1.602 * 10-19) / (1.673 * 10-27) = 5.65 * 1010 m/s2

(b.) v = u+at

1.3 * 106 = 5.65 * 1010 * t

t = 0.23 * 10-4 s

(c.) S = v2 / 2a = 1.69 * 1012 / (2 * 5.65 * 1010 ) = 14.956 m = 15m

(d.) K.E = 0.5 mv2 = 0.5 * 1.673 * 10-27 * 1.69 * 1012 = 1.41 * 10-15 J
Answer #2

a.E=F/q

Eq=F=ma


a= Eq/m , you know the mass and charge of proton and E is given

b. use the newton equation . intitial velocity is zero and final is given. so its S=vit + 1/2 at^2

c) 2as=vf^2-vi^2

d) ke=1/2 mv^2

now i think you are able to put the value and solve it.